Supplemental Operation – February 25th

Supplemental Operations Report 25Feb2022

Name of person filing report: Shannon Rupert

Reason for Report: Changing of the guard

Non-nominal systems: GreenHab

Action taken for non-nominal systems: All crops were removed from the GreenHab and it will be shut down quite possibly for the rest of the field season

Generator: Still being manually run at night. 967.5 hours

ScienceDome Dual Split: Off

Solar—Charge controllers off at night. We turn the generator on before the SOC reaches 70%.

Solar— VDC Last daylight hours: (information obtained from magnum data)

Average 52.31

Minimum 48.40

Maximum 59.20

Notes on solar: Nothing to report

Propane Reading, station tank – 70 %

Propane Reading, director tank— 65 %

Propane Reading, intern tank— 55 %

Propane Reading, generator— 55 %

Ethanol Free Gasoline – 2 gallons

Water (Outpost tank) – 200 gallons

Sojourner rover used: Yes

Hours: 161.9

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): No

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, local

CrewCar used and why, where? Yes. Atila took it to drive to SLC to pick me up but lost control and went off the road close to I-70 on Highway 24. He was shaken and sore afterwards but the car was checked and appears to have no damage.

Luna used and why, where? No.

Campus wide inspection, if action taken, what and why? I have not completed my inspection fully but noted no damage/changes to the Robotic Observatory exterior, instrumentation nearby or the solar panels. Signs on the road and at the entrance have no damage/changes.

General notes and comments: I arrived back at MDRS on Wednesday night and Atila will be leaving MDRS on Saturday morning.

Summary of general operations:

Summary of internet: The original Hab internet service will be deactivated on March 1, 2022. Today the new system was to be installed, but the installer cancelled. I will be installing a Starlink system that hopefully can be used while we are waiting for HughesNet to get its act together.

Summary of suits and radios: Nothing to report

Summary of Hab operations: The toilet holding tank was cleaned for the built up waste that was making the tank fill twice as fast as when nominal. We were unable to see how successful we were, but will track the crew’s use to see if we cleared it completely. We will repeat this process in 4 weeks.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I inspected the GreenHab the morning after I arrived. From what few details I was getting from reports and photos, it was clear something was wrong. Nothing was thriving and plants were dying and underproducting. A more detailed inspection showed an unchecked infestation of white fly and contamination of soil by eggs and larva. I determined a collapse was near, salvage wasn’t possible, and we shut the building down. The crew removed all crops and consolidated the soil, and I will be finishing the removal of plant debris over the next few days. Heat is off and supplemental lighting turned off. This failure was the result of not one, but many small things. I was offsite and hadn’t been paying enough attention to the reports to realize that the crew was not reporting vital information about the condition of the plants. Our supplemental operations report had become truncated in a way that didn’t allow clear reporting on GreenHab issues by Atila. The GreenHab Officers and Atila didn’t have the knowledge and experience to recognize what was happening and I was offsite. Although we are requiring a general skill set for GreenHab Officer’s beginning next season, we do not have any requirements at the moment and it is one of two positions that are generally not filled by qualified people. Finally, but maybe not of much importance, I requested that the crew install a shade cloth weeks ago, and that was never gone. Dying plants were reported as drying plants. This is, however, why we do what we do—to learn from what mistakes we make. We probably won’t be able to produce anymore crops this season, because we need to sterilize the soil before we can replant and that requires heat, sunlight and time.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Robotic observatory is down. A part to repair it is due here Saturday.

Summary of health and safety issues: Nothing to report

Journalist Report – February 25th

Sol 4 – Experiments filled day

Author : Nicolas WATTELLE

To be honest, today I felt a bit tired. The accumulation of the previous days and the double EVAs of the morning might be responsible for that.

First, two EVAs were planned during the day. One during the morning, one for the afternoon. But because of the snow melting, we had to condense the two outings in half a day: the first team was made by Matheo, Elena, Valentine and I. For the second, we just went out with Matheo. We tried for the first time, Matheo’s Search and Rescue protocol with drones. For him it was a bit stressful, because he was worrying for the possible unforeseen of piloting a drone in simulation conditions. While for Elena and I it was way more relaxing: we just had to walk in a predefined zone, and… “get lost”. The trial went well, Valentine and Matheo geotagged us very quickly, and we left the zone with some ideas to improve the protocol; it is promising.

In the afternoon, we were mostly all out the hab. Some of us were working on repairing the MegaAres antenna. It is interesting to tinker in a constrained area, with limited resources: “do what you can with what you have” is a sentence that seems like a chorus for Mathéo and I. Until now, we always managed to go further in this tricky repair, let’s hope it lasts! Valentine, Marine and Lea performed sessions of some of our human factor experiments: Teleop and other tests grouped in a software called “TAP”. TAP is definitively our daily nightmare. Briefly, imagine that you have to stay in front of your computer during an hour and a half, answering questions and doing cognitive exercises during which you test your focus in various ways. Sometimes you need to find correlations between shapes, colours, sounds… and “tap” on the right button as soon as you can. The first time it felt like challenging, almost like a game, but once you have done it several times it becomes really boring… But we are dedicated to our mission and we believe these data can help researchers. At the end of the day, it remains a subject that makes us laugh during meals so fair enough!

Today, in my tired person thoughts, I asked my teammates: How do you imagine the first crew walking on Mars?

That started a debate, even if everyone agreed on most subjects. Léa talked about the importance to send experts in widely various domains: engineering, medicine, geology… Marine and others hope that women and men would be equally represented, and more generally that it would be a fair selection. One question that was raised, is the one of the nationalities of the crew members. Does a country will succeed doing it alone, or would it be an international collaboration? Here at the MDRS, some are really convinced, and some are not. Only the future will tell us!

Scientific Report – February 25th

Crew 263 Scientific Report 25-02-2022

SOL: 4

Name of person filing report: Léa Rouverand

Ultrasound surveillance:

On SOL 2, image capture of 5 organs for each Crew Members were performed using an ultrasound device and an Augmented Reality software. No issues were found regarding the hardware. Some issues in the protocol were identified and sent to experts at CNES (French Space Agency).

HF wave propagation experiment:

This experiment aims to study the propagation characteristics of the ionosphere for HF waves to achieve contact between MDRS and Toulouse (France). The Delta Loop Antennas were successfully set up on SOL 3.


Rotating platforms to study the effect of gravity on plant growth which were initially installed in the GreenHab were moved to the Science Dome on SOL 3.

Drone Astronaut Rescue:

The experiment’s aim is to assess the use of an ANAFI thermal drone for Astronaut Rescue. First trial was performed on SOL 4, the drone allowed to spot the Astronaut pair in 2 minutes and join them in 26 minutes. The use of the drone thermal feature is still to be investigated as it made the image more confusing. We also need to compare the rescue protocol to one without a drone; this is planned on SOL 16.

Atmospheric Experiment:

There is a total of 5 atmospheric experiments; LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a weather station. On SOL 4, the installation of 4 of them was finished. Mega-Ares still needs to be installed.

Augmented Reality Testing:

First trials of the Augmented Reality Application testing were performed between SOL 1 and SOL 4. Software issues with menu placement and hardware issues were identified.

On-going experiments: Sleep study using Dreem Headbands, TELEOP (analysis of spatial awareness and isolation on teleoperation tasks), Psychosocial and Cognitive questionnaires, Psychometric tests, Water quality monitoring. Nothing to declare on SOL 4.

Questions, concerns and requests to Mission Support: NaN

Operations Report – February 25th

Crew 263 Operations Report 25-02-2022

SOL: 5

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

Opportunity rover used: Yes

Hours: 84.7

Beginning charge: 100%

Ending charge: 77%

Currently charging: left uncharged, handled by Atila and Shannon

Curiosity rover used: Yes

Hours: 168.7

Beginning charge: 100%

Ending charge: 76%

Currently charging: left uncharged, handled by Atila and Shannon

Perseverance rover used: No

Hours: 84.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila and Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 15 gallons

Water (static tank): 443 gallons

Water (loft tank): 40 gallons

Water Meter: 01561172 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: NA

Summary of suits and radios: NA

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: 74% at 6pm

Summary of RAM operations: We used the drill, the soldering iron and the electric fill to repair an atmospheric experiment named MegaAres that measure the conductivity of particles in the air

Summary of any observatory issues: There is an issue with the dome opening, the first part of the dome opens perfectly but the second part is not moving anymore. Yesterday and the day before, the whole dome was perfectly functioning.

Summary of health and safety issues: See HSO report

Questions, concerns and requests to Mission Support: We put trash in the rear airlock for pickup. Thanks a lot!

EVA Report_2 – February 25th

Crew 263 EVA Report_2 25-02-2022

EVA # 6

Author: Nicolas WATTELLE

Purpose of EVA: Testing the Hololens waypoints protocol

Start time: 11:30

End time: 12:10

Narrative: Nicolas was wearing the Hololens. The device had some trouble working at the beginning of the outing so they left the airlock late. Mathéo and Nicolas walked around and stopped in front of every room of the station. At each stop, Nicolas had to place a waypoint with the Hololens and take photos. Then, they walked back on each point to check if the waypoints were still there.

Destination: Around the station.

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants: Nicolas Wattelle, Mathéo Fouchet

Road(s) and routes per MDRS Map: NA

Mode of travel: foot.

EVA Report_1 – February 25th

Crew 263 EVA Report_1 25-02-2022

EVA # 5

Author: Mathéo Fouchet

Purpose of EVA: Augmented Reality application testing to map MDRS, Battery replacement for atmospheric experiment, Search and Rescue exercise with Drone within a 1 mile radius.

Start time: 8:08

End time: 10:01

Narrative: We checked the rover batteries (all full), and the Red team left with Opportunity to reach our atmospheric experiment. During this time the Blue team was mapping the station with Hololens. There was a problem with this experiment because we were not able to see the way points put with the Hololens. In addition, we forgot a cable to plug the LOAC (Light Optical Aerosol Counter) on the battery. Thus, when the Blue team finished the Augmented Reality experiment , they had to wait for the depressurisation of the airlock and got the missing cable at 8:30. The Blue team then joined the Red team at the atmospheric experiment with Curiosity at 8:34. All the Atmospheric experiments were then set up and working correctly.

To implement the Search and Rescue with Drone experiment, the Red team left at 8:39 with Opportunity (97% battery) and stopped at 8:42 (87% battery) near North Ridge while the Blue team was waiting at the atmospheric experiment. Then, the Red team walked to reach the north of North Ridge. At 8:56, the Blue team took the Curiosity Rover (98% battery) and stopped behind the Red team’s rover at 9:04 (86% battery for curiosity at this time). The drone was launched at 9:08 and the thermal camera was turned on in order to look for the Red team. The Red team was spotted only 2 minutes after the launch. The Blue team then used the drone to find the easiest way to reach the Red team and both teams gathered at 9:29.

We finally reached the rovers at 9:48 and reached the Hab at 10:01.

Destination: Around the Hab, Intersection between Entrance Road and Cow Dung Road, and North Ridge

Coordinates (use UTM NAD27 CONUS): 518250 E, 4250750 N; 518288E, 4250586N; 518500E, 4251800N

Participants: Valentine Bourgeois,Mathéo Fouchet, Nicolas Wattelle, Elena Lopez-Contreras.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road.

Mode of travel: Foot, Rover

Astronomy Report – February 25th

Name: Marine Prunier

Crew: 263
Date: 25 Feb 2022


Robotic Telescope Requested (choose one): MLC-RCOS16

Objects to be Imaged this Evening: None

Images submitted with this report: M81

Problems Encountered: /


Solar Features Observed: /

Images submitted with this report: /

Problems Encountered: The second part of the dome is not opening anymore. Yesterday the whole dome was perfectly functioning.

Sol Summary – February 25th

Crew 263 Sol Summary Report 25-02-2022

Sol: 4

Summary Title: Productive work day.

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: While four of us were on EVA, the others (including me) did a sport session, swept the lower Hab and started the Mars puzzle. For lunch we cooked a delicious french gratin dauphinois -it is good to know that it is possible to do béchamel with dehydrated butter! In the afternoon, some of us were occupied with prepping MegaAres -that we might finally setup outside with the rest of the atmospheric experiments by the beginning of next week, Human Factors experiments that were exhausting, HF communication tests, and astronomy post-processing. Some bread is in the oven for tonights diner.

Look Ahead Plan: We are looking forward to trying communication with Toulouse tomorrow with our HF antenna. Hopefully the weather and terrain conditions will be good enough to do our first LIBS analysis during EVA.

Anomalies in work: None.

Weather: Mostly sunny.

Crew Physical Status: Good.

EVA: Two EVAs in the morning, both were successful.

Reports to be filed: Sol Summary, 2 EVA Requests, 2 EVA Reports, Operations Report, Journalist Report, HSO Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site